# Entangling Quantum Memories at Channel Capacity

### v1

@@ Abstract
Entangling quantum memories, mediated by optical-frequency or microwave channels, at high rates and fidelities is key for linking qubits across short and long ranges. All well-known protocols encode up to one qubit per optical mode, hence entangling one pair of memory qubits per transmitted mode over the channel, with probability $\eta$, the channel’s transmissivity. The rate is proportional to $\eta$ ideal Bell states (ebits) per mode. The quantum capacity, $C(\eta)=-\log_{2}(1-{\eta})$ ebits per mode, which $\approx 1.44\eta$ for high loss, i.e., $\eta\ll 1$, thereby making these schemes near rate-optimal. However, $C(\eta)\to\infty$ as $\eta\to 1$, making the known schemes highly rate-suboptimal for shorter ranges. We show that a cavity-assisted memory-photon interface can be used to entangle matter memories with Gottesman-Kitaev-Preskill (GKP) photonic qudits, which along with dual-homodyne entanglement swaps that retain analog information, enables entangling memories at capacity-approaching rates at low loss. We benefit from loss resilience of GKP qudits, and their ability to encode multiple qubits in one mode. Our memory-photon interface further supports the preparation of needed ancilla GKP qudits. We expect our result to spur research in low-loss high-cooperativity cavity-coupled qubits with high-efficiency optical coupling, and demonstrations of high-rate short-range quantum links.
@@
Entanglement is a key resource for all tasks in quantum information processing, that promises an advantage over conventional classical techniques. The task of generating entanglement between multiple parties is therefore a core challenge, be it among multiple qubits on a quantum processor, between many nodes on a metropolitan scale quantum network, or over long haul networks connecting quantum data centers relying on quantum repeaters, aerial nodes or satellite links. The fundamental unit of any such network is a single quantum link, connecting two parties over a lossy optical (or potentially microwave) frequency channel, most often simply characterized by the channel transmissivity $\eta$.
@@
Most prevalent quantum link protocols rely on some encoding of the quantum information in the optical mode that encodes up to one qubit per optical mode transmitted – limiting them to generate entanglement between at most a pair of quantum memories (ideally, in one of four Bell states). This may succeed at a rate proportional to $\eta$ ideal Bell states (ebits) per mode. The maximum allowable entanglement rate is the quantum capacity of the pure loss channel, $C(\eta)=-\log(1-{\eta})$ ebits per mode. This scales roughly as $\approx 1.44\eta$ for high loss, i.e., $\eta\ll 1$, thereby making the aforesaid schemes near rate-optimal. However, $C(\eta)\to\infty$ as $\eta\to 1$. Therefore, for shorter ranges, viz., intra-processor, data-center, local-area quantum networks and potentially even long haul networks relying on ultra-low loss transmission assisted by vacuum beam guides, the known schemes are highly rate-suboptimal.
In our article, we propose a rate-optimal quantum link for operating in the low loss regime. We show that a cavity-assisted memory-photon interface [1] can be used to entangle $N$-matter quantum memories with higher dimensional bosonic qudit encoding i.e. encoding $d>1$ qubits per optical mode, where $d=2^{N}$. We utilize qudit subspaces of the infinite dimensional bosonic mode, whose logical basis states are defined by their orthonormal phase space distributions. We use the Gottesman-Kitaev-Preskill (GKP) photonic qudits [2] for the memory-photon entanglement step, to generate (in the ideal scenario) a maximally entangled $d$-dimensional Bell pair, which is the key resource state. Performing dual-homodyne-based entanglement swaps (that retain the analog information [3, 4]) on the bosonic modes of two resource states, enables entanglement of the pair of $N$ memories at a rate approaching capacity at low loss.
Our scheme benefits from the resilience of GKP qudits to channel loss [5, 6] and random displacements [2]. Furthermore, the GKP encoding can be naturally extended to accomodate multiple qubits in one mode, which is not possible with the more-common single-rail, dual-rail, and multi-mode high-dimensional single-photon qubit encodings. Additionally, it has been shown that a similar memory-photon interface naturally supports the preparation of ancilla GKP qubits our scheme needs [7, 8].
@@
Our article is structured as follows – we briefly introduce the GKP encoding and the underlying assumption for our quantum memories. We then describe the memory photon interface to generate the $d$-dimensional Bell pairs, relying on an effective CSUM gate protocol. Finally, we demonstrate the rate-optimality of the proposed quantum link in the low-loss regime. We also propose potential applications of the CSUM gate protocol.
Our results show that even with $5$-dB-squeezed GKP qubits, one could double intra-processor entanglement rates compared with the theoretical limits of known schemes. We anticipate our result to lead to more sophisticated protocols and practical demonstrations of high-rate quantum links. Furthermore, developing low-loss high-cooperativity cavity-coupled qubits with high-efficiency optical coupling is a major target, and our scheme will further catalyze this technology development.
@@
We consider a single bosonic mode with the creation ($\hat{a}$) and annihilation ($\hat{a}^{\dagger}$) operators defined in terms of the canonical position ($\hat{q}$) and momentum ($\hat{p}$) as $\hat{a}=(\hat{q}+i\hat{p})/\sqrt{2};\hat{a}^{\dagger}=(\hat{q}-i\hat{p})/\sqrt{2}$ (i.e. $\hbar=1$ units). The logical square-lattice GKP (sq-GKP) qudit subspace [2, 6] is defined as the eigenspace formed by the displacement operators
$\begin{split}\hat{S}_{q}^{(d)}&=\exp{(-i\sqrt{2\pi d}\hat{p})};\\ \hat{S}_{p}^{(d)}&=\exp(i\sqrt{2\pi d}\hat{q}).\end{split}$ (1)
@@
where $\mu_{q}=0,1,\ldots,d-1$. It is straightforward to note that a position quadrature displacement of $j\sqrt{2\pi/d}$, maps states from |$\ket{\mu_{q}}\rightarrow\ket{\mu_{q}+j}$. Note that any logical state label $\mu_{q}$ is an integer in the residue set of the qudit dimension $d$; for the rest of article we will drop an implicit modulo in our description of the states. Another natural choice is the hexagonal-lattice GKP (hex-GKP) qudit subspace, which is defined as the eigenspace of the displacement operators
$\begin{split}\hat{S}_{q}^{(d)}&=\exp{[i\sqrt{2\pi d}\left({2}/{\sqrt{3}}\right)^{1/2}(\hat{q}+\hat{p}/2)]};\\ \hat{S}_{p}^{(d)}&=\exp{[-i\sqrt{2\pi d}\left({2}/{\sqrt{3}}\right)^{1/2}(\sqrt{3}\hat{p}/2)]}.\end{split}$ (3)
As with all GKP encodings [2], the ideal code states defined above are not physical owing to their non-normalizable definition (i.e. infinite superpositions of quadrature eigenstates); physical approximations to the ideal states can be made by imposing finite peak widths and/or phase envelopes. For the rest of the article, we shall consider randomly displaced versions of the ideal logical states, defined as
@@
where $\Gamma(u,v)$ is a bi-variate Gaussian displacement function,
$\Gamma(u,v)=\frac{1}{2\pi{\sigma_{u}\sigma_{v}}}\exp\left[-\frac{u^{2}}{2\sigma_{u}^{2}}-\frac{v^{2}}{2\sigma_{v}^{2}}\right].$ (5)
Although this definition is still un-physical (i.e. due to the infinite superposition), the effect of physical operations and channels on the qudit states can be effectively treated as a evolution of the random displacement variances (i.e. $\sigma^{2}_{u},\sigma^{2}_{v}$).
@@
Our protocol relies on a cavity-assisted interaction of multiple bosonic coherent pulses with atomic systems to induce the required optical non-Gaussian operations. We consider $N$ three-level atomic systems (which we shall henceforth refer to as the ‘memories’) each with a single coupled interaction at optical frequencies. Levels |$\ket{0}_{M}$ and |$\ket{1}_{M}$ form the logical qubit - additionally the |$\ket{1}_{M}$ level is optically coupled to a third level, |$\ket{2}_{M}$. A variety of physical qubit platforms can support such a level structure - charged quantum dots [9, 10], trapped ions [11], and solid state vacancy centers [12, 13, 14]. Each atomic system is integrated in a single-mode cavity with one coupling port that couples the cavity mode ($\hat{a}_{c}$) with incoming ($\hat{a}_{\mathrm{in}}$) and outgoing ($\hat{a}_{\mathrm{out}}$) free-space modes at coupling rate $\kappa_{c}$ (depicted in Fig. 1). With the memory system initialized in its qubit subspace, the incoming optical field is reflected with a memory state dependent phase, effectively implementing an unitary transformation $C\!R_{\pi}$ [15, 16], where,
@@
Figure 2: Overview of the interaction of the atomic system with the GKP qubit for the effective CSUM gate – (a) Overview of the protocol implementing the CSUM gate, involving reflection of coherent pulses from a register of $N$-quantum memories integrated in cavities and memory-state dependent sequential displacement of pre-dislpaced GKP qudit states. (b) Circuit description of the protocol comprising of sequential controlled phase rotations $(\hat{U}_{\mathrm{R}})$ and sequential displacements $(\hat{U}_{\mathrm{disp.}})$.
@@
Without loss of generality, consider the case where the initial state of the GKP qudit is |$\ket{0}_{G}$ (with $d=2^{N}$). The state is pre-displaced to the |$\ket{\phi}_{G}=D((d-1)\sqrt{2\pi/d}/2)\ket{0}_{G}$ state. Simultaneously, the memory register is initialized in the |$\ket{\psi}_{\bm{M}}=\bigotimes_{i=1}^{N}\ket{+}_{M_{i}}$ state. This can equivalently be expressed as
@@
    Reflect a coherent pulse |$\ket{\alpha_{k}}_{B_{k}}$ in the $B_{k}$ mode from the memory $M_{k}$. The coherent amplitudes are specified by,
@@
The exact phase rotation of the reflected bosonic pulse is a critical factor for the CSUM gate performance –any imperfection in the interface limit sthe effectiveness of the protocol. This shows up as an effective dephasing noise on the final state. We have explored the efficacy of a non-ideal protocol implementation for $N=1;d=2$ protocol in detail in Ref. [1]; here we highlight the main limitations posed by these interactions.
Assuming the same state initialization as described previously for the protocol, the final state can most generally be described by,
$\begin{split}\rho_{\bm{M}G}\approx\sum_{\bm{m},\bm{m}^{\prime}=0}^{d-1}g_{\bm{m},\bm{m}^{\prime}}\ket{\bm{m}}\!\!\bra{\bm{m}^{\prime}}_{\bm{M}}D(\bm{\beta}_{\bm{m}^{\prime}})\rho_{G}D^{\dagger}(\bm{\beta}_{\bm{m}^{\prime}}),\end{split}$ (11)
where corresponding to the memory basis state |$\ket{\bm{m}}_{\bm{M}}$ the realistic displacement is
@@
$g_{\bm{m},\bm{m}^{\prime}}=\left(\delta_{\bm{m},\bm{m}^{\prime}}+(1-\delta_{\bm{m},\bm{m}^{\prime}})|\lambda_{\bm{m},\bm{m}^{\prime}}|^{2}\right)/d,$ (13)
with $|\lambda_{\bm{m},\bm{m}^{\prime}}|^{2}$ representing the effective dephasing factors governed by photons lost by imperfect cavity coupling and atomic decay. For $C\gg 1$ and $|\omega-\omega_{c}|\rightarrow 0$, we may achieve the approximate relation $-r^{(0)*}(\omega)\approx r^{(1)*}(\omega)=1$ over the spectral bandwidth of the pulses (see Methods for details). Any additional sources of loss can be incorporated by including sub-unity multiplicative factors for $|\lambda_{\bm{m},\bm{m}^{\prime}}|^{2}$.
@@
Figure 3: Layout of high-rate quantum links utilizing GKP qudit basis entanglement swap operation, where each party performs the CSUM gate on local memory registers $(\bm{M}_{1},\bm{M}_{2})$ and GKP qudit states $(G_{1},G_{2})$ before transmission to the midpoint of a link (of transmissivity $\eta$).
@@
The effect of loss (with implicit pre-amplification) and finite-squeezing on the GKP qudit states can be effectively modeled by a Gaussian random displacement channel on the ideal qudit states [4, 18, 6]. Given an initial state described by the density operator $\rho_{G}$, the random displacement channel $\mathcal{N}_{B_{2}}[\sigma^{2}]$ modifies the state as
@@
Considering the case $\eta=1$ (i.e. lossless transmission), the final state is |$\ket{\Psi_{k^{\prime},l^{\prime}}}\equiv W_{1}^{(0,d-x_{L})}W_{2}^{(d-y_{L},0)}\ket{\Psi_{0,0}}$ given the Bell state measurement logical outcomes $x_{L},y_{L}$. |$\ket{\Psi_{k^{\prime},l^{\prime}}}$ is a different qudit Bell state; since both $x_{L}$ and $y_{L}$ have $d$ possible values, the heralded state is one of $d^{2}$ qudit Bell states (see Methods for notation related to qudit gates).
Assuming that the parties have access to finitely-squeezed GKP states of peak variance $\sigma_{i}^{2}$, transmission over a pure loss channel of transmissivity $\eta$ can be treated as a modification of the peak variance by $\sigma_{i}^{2}\rightarrow\sigma_{i}^{2}+(1-\eta)$. Given a state variance of $\sigma^{2}$, the probability of making $k$-shift errors in the logical measurement for sq-GKP states (along either the Pauli $X$ or $Z$ basis) is given by [6, 19],
$\begin{split}&P^{(\mathrm{sq})}(X^{k},\sigma^{2})\equiv P^{(\mathrm{sq})}(Z^{k},\sigma^{2})\\ &=\sum_{j\in\mathbb{Z}}\frac{1}{2}\left(\mathrm{erf}\left(\sqrt{\frac{2\pi}{d}}\frac{jd+k+1/2}{\sigma}\right)-\mathrm{erf}\left(\sqrt{\frac{2\pi}{d}}\frac{jd+k-1/2}{\sigma}\right)\right),\end{split}$ (18)
where $k\in\{0,1,2,\ldots,d-1\}$. $k=0$ denotes the probability of incurring no shifts; the summation over the index $j$ accounts for shifts of magnitude greater than one lattice spacing of $\sqrt{2\pi d}$. Assuming that Bell state measurement logical outcomes are $x_{L},y_{L}$ the final state can be expressed as
$\tilde{\rho}_{\bm{M}_{1}\bm{M}_{2}}=\sum_{k_{1},k_{2}=0}^{d-1}P_{\mathrm{shift}}^{(k_{1},k_{2})}(\sigma^{2})\,\ket{\Psi_{k^{\prime},l^{\prime}}}\!\!\bra{\Psi_{k^{\prime},l^{\prime}}}_{\bm{M}_{1}\bm{M}_{2}}$ (19)
where $P_{\mathrm{shift}}^{(k_{1},k_{2})}(\sigma^{2})=P^{(\mathrm{sq})}(X^{k_{1}},\sigma^{2})\cdot P^{(\mathrm{sq})}(Z^{k_{2}},\sigma^{2})$. Assuming a mixed state model for the final state, the hashing bound based lower bound to the distillable entanglement is given by

### v2

@@ Abstract
Entangling quantum memories, mediated by optical-frequency or microwave channels, at high rates and fidelities is key for linking qubits across short and long ranges. All well-known protocols encode up to one qubit per optical mode, hence entangling one pair of memory qubits per transmitted mode over the channel, with probability $\eta$, the channel’s transmissivity. The rate is proportional to $\eta$ ideal Bell states (ebits) per mode. The quantum capacity, $C(\eta)=-\log_{2}(1-{\eta})$ ebits per mode, which $\approx 1.44\eta$ for high loss, i.e., $\eta\ll 1$, thereby making these schemes near rate-optimal. However, $C(\eta)\to\infty$ as $\eta\to 1$, making the known schemes highly rate-suboptimal for shorter ranges. We propose a cavity-assisted memory-photon interface that can be used to entangle matter memories with Gottesman-Kitaev-Preskill (GKP) photonic qudits, which along with dual-homodyne entanglement swaps that retain analog information, enables entangling memories at capacity-approaching rates at low loss. We benefit from loss resilience of GKP qudits, and their ability to encode multiple qubits in one mode. Our memory-photon interface further supports the preparation of needed ancilla GKP qudits. We expect our result to spur research in low-loss high-cooperativity cavity-coupled qubits with high-efficiency optical coupling, and demonstrations of high-rate short-range quantum links.
@@
Entanglement is a key enabling resource for many tasks in quantum enhanced information processing. The task of generating entanglement at high rates and high fidelity between multiple parties is a core challenge, be it among multiple qubits in a quantum processor, between many nodes of a metropolitan-scale quantum network, or over long haul networks connecting quantum data centers relying on quantum repeaters, aerial nodes or satellite links. The fundamental unit of any such network is a single quantum link, connecting two parties over a lossy optical (or potentially microwave) frequency channel, often characterized by the channel’s transmissivity $\eta$.
@@
Most prevalent protocols for generating heralded entanglement among qubits in quantum memories connected by a lossy channel rely on quantum-optical modulation formats that encode up to one qubit per transmitted optical mode. This limits each transmitted mode to attempt to establish entanglement between at most a pair of distant qubits. If this attempt succeeds with probability proportional to $\eta$, e.g., when the dual-rail encoding is used which encodes one qubit using one photon spread across two modes, then the rate at which entanglement is generated is proportional to $\eta$ ideal Bell states (ebits) per mode. The maximum allowable entanglement rate, assisted with two-way authenticated classical communications, is the quantum capacity of the pure loss bosonic channel, given by: $C(\eta)=-\log(1-{\eta})$ ebits per mode. This scales roughly as $\approx 1.44\eta$ in the high loss, i.e., $\eta\ll 1$ regime, thereby making the known schemes near rate-optimal in the high-loss regime. However, $C(\eta)\to\infty$ as $\eta\to 1$. Therefore, for shorter ranges, viz., intra-processor and data-center quantum links, local-area quantum networks and potentially even long haul networks relying on ultra-low loss transmission assisted by vacuum beam guides [1], the known schemes are highly rate-suboptimal.
In this Article, we show that a cavity-assisted memory-photon interface we recently proposed [2] can be used to entangle $N$-qubit blocks of two remote memory registers, mediated by $d$-dimensional photonic (qudit) encodings at rates approaching capacity at low channel loss [2]. Our proposal uses individual devices that have been demonstrated in isolation, hence making it an attractive candidate for near-future realizations. We first generate – at each of the two sites – a maximally-entangled memory-photon entangled state of an $N$-qubit memory register and a single optical mode encoding $N=\log_{2}(d)$ qubits in the Gottesman-Kitaev-Preskill (GKP) photonic qu(d)it format [3]. The optical qudits from each site undergo lossy propagation to a middle station where a dual-homodyne-based entanglement swap is performed, whose analog outcome is retained for further post-processing of the memory-memory remote entanglement created, at the error correction stage [4, 5]. If the channels have zero loss, and for an ideal cavity-assisted memory-photon interface, our protocol generates $N$ ebits per transmitted mode where $N$ can be arbitrarily high in principle, in congruence with the fact that $C(\eta)=\infty$ when $\eta=1$. In the low-loss regime ($\eta$ approaching $1$), we show that the distillable entanglement rate achieved by our protocol approaches the quantum capacity, and outperforms known schemes by order of magnitude or more, even with finite-squeezed GKP qudits.
Our scheme benefits from the resilience of GKP qudits to channel loss [6, 7] which manifests as random Gaussian displacements in the phase space for GKP qudits [3]. Furthermore, the GKP encoding can be naturally extended to accommodate multiple qubits in one mode, which is not possible with the more-common single-rail, dual-rail, cat-basis and single-photon multi-mode high-dimensional qubit encodings. It has also been shown that a similar memory-photon interface naturally supports the preparation of ancilla GKP qubits our scheme needs [8, 9]. Our results show that even with $5$-dB-squeezed GKP qubits, one could double intra-processor entanglement rates compared with the theoretical limits of known schemes utilizing single photon based photonic qubit encoding and linear optical entanglement swaps. Furthermore, by choosing an optimal encoding size $d$ commensurate to the channel loss, our protocol approaches capacity within a constant separation of $\sim\!1$ ebit per channel use. We anticipate our result to lead to more sophisticated protocols and practical demonstrations of high-rate short-range quantum links. Extensions of our protocol are applicable to a variety of research questions including hybrid continuous-discrete variable quantum logic gates and resource efficient non Gaussian photonic state preparation. Furthermore, developing low-loss high-cooperativity cavity-coupled qubits with high-efficiency optical coupling is a major experimental target, and our scheme will further catalyze this technology development.
@@
Our article is structured as follows. We briefly introduce the GKP encoding and the underlying assumptions on our atomic quantum memories. We then describe the memory-photon interactions required to generate the spin-photon $d$-dimensional Bell pairs; implementing an effective CSUM gate. Finally, we evaluate the performance of the proposed quantum links while demonstrating their rate-optimality in the low-loss regime, and propose other potential applications of the CSUM protocol.
@@
We consider a single bosonic mode with the creation ($\hat{a}$) and annihilation ($\hat{a}^{\dagger}$) operators defined in terms of the canonical position ($\hat{q}$) and momentum ($\hat{p}$) as $\hat{a}=(\hat{q}+i\hat{p})/\sqrt{2};\hat{a}^{\dagger}=(\hat{q}-i\hat{p})/\sqrt{2}$ (i.e. $\hbar=1$ units). The logical square-lattice GKP (sq-GKP) qudit subspace [3, 7] is the eigenspace formed by the displacement operators:
$\begin{split}\hat{S}_{q}^{(d)}&=\exp{(-i\sqrt{2\pi d}\hat{p})},\,{\text{and}}\\ \hat{S}_{p}^{(d)}&=\exp(i\sqrt{2\pi d}\hat{q}).\end{split}$ (1)
@@
where $\mu_{q}=0,1,\ldots,d-1$. It is straightforward to note that a position quadrature displacement of $j\sqrt{2\pi/d}$, maps states from |$\ket{\mu_{q}}\rightarrow\ket{\mu_{q}+j}$. Note that any logical state label $\mu_{q}$ is an integer in the residue set of the qudit dimension $d$; for the rest of article we will drop an implicit modulo in our description of the states. Another natural choice is the hexagonal-lattice GKP (hex-GKP) qudit subspace, which is defined as the eigenspace of the displacement operators:
$\begin{split}\hat{S}_{q}^{(d)}&=\exp{[i\sqrt{2\pi d}\left({2}/{\sqrt{3}}\right)^{1/2}(\hat{q}+\hat{p}/2)]},\,{\text{and}}\\ \hat{S}_{p}^{(d)}&=\exp{[-i\sqrt{2\pi d}\left({2}/{\sqrt{3}}\right)^{1/2}(\sqrt{3}\hat{p}/2)]}.\end{split}$ (3)
As with all GKP encodings [3], the ideal code states defined above are not physical owing to their non-normalizable definition, i.e., infinite superpositions of quadrature eigenstates. Physical approximations to the ideal states can be made by imposing finite peak widths and/or phase envelopes. In this Article, we will take finite-squeezed physical approximations as random-Gaussian displacements of the ideal logical states:
@@
where $\Gamma(u,v)$ is a bi-variate Gaussian distribution,
$\Gamma(u,v)=\frac{1}{2\pi{\sigma_{u}\sigma_{v}}}\exp\left[-\frac{u^{2}}{2\sigma_{u}^{2}}-\frac{v^{2}}{2\sigma_{v}^{2}}\right],$ (5)
where typically $\sigma_{u}^{2}=\sigma_{v}^{2}=\sigma^{2}$. Although this definition is still not entirely practical (i.e., due to the infinite superposition in (2)), the effect of channel loss and noise on GKP qudits can be effectively treated as evolution of the Gaussian-displacement variance $\sigma^{2}$, hence making it analytically tractable.
@@
We consider $N$ three-level atomic systems (henceforth referred to as the memories) each with a single coupled interaction at an optical frequency (see Fig. 1). Levels |$\ket{0}_{M}$ and |$\ket{1}_{M}$ form the logical qubit. The |$\ket{1}_{M}$ level is optically coupled to a third level, |$\ket{2}_{M}$. A variety of physical qubit platforms can support such a level structure, e.g., charged quantum dots [10, 11], trapped ions [12], cold atoms [13], and solid state vacancy centers [14, 15, 16]. Each atomic system is integrated in a single-mode cavity with one coupling port that couples the cavity mode ($\hat{a}_{c}$) with incoming ($\hat{a}_{\mathrm{in}}$) and outgoing ($\hat{a}_{\mathrm{out}}$) free-space optical modes at coupling rate $\kappa_{c}$. With the memory system initialized in its qubit subspace, the incoming optical field is reflected with a memory-state-dependent phase, effectively implementing an unitary transformation $CR_{\pi}$ [13, 17], where,
@@
Figure 2: Overview of the interaction of the atomic system with the GKP qubit for the effective CSUM gate – (a) Overview of the protocol implementing the CSUM gate, involving reflection of coherent pulses from a register of $N$-quantum memories integrated in cavities and memory-state dependent sequential displacement of pre-displaced GKP qudit states. (b) Circuit description of the protocol comprising of sequential controlled phase rotations $(\hat{U}_{\mathrm{R}})$ and sequential displacements $(\hat{U}_{\mathrm{disp.}})$.
@@
Without loss of generality, consider the case where the initial state of the GKP qudit is |$\ket{0}_{G}$ (with $d=2^{N}$). The state is pre-displaced to the |$\ket{\phi}_{G}=D((d-1)\sqrt{2\pi/d}/2)\ket{0}_{G}$ state, where $D(\alpha)=\exp(\alpha\hat{a}^{\dagger}_{G}-\alpha^{*}\hat{a}_{G})$ is the displacement operator for the bosonic mode excited in the GKP qudit state. Simultaneously, the memory register is initialized in the |$\ket{\psi}_{\bm{M}}=\bigotimes_{i=1}^{N}\ket{+}_{M_{i}}$ state. This can equivalently be expressed as
@@
    Reflect a pulse in the coherent-state |$\ket{\alpha_{k}}_{B_{k}}$ of the $B_{k}$ mode off of the cavity-coupled memory $M_{k}$. The coherent amplitudes are specified by,
@@
Any imperfection in the cavity photon interface limits the effectiveness of the proposed protocol. Primarily, the cavity-photon interface must impart the ideal memory-dependent phase while minimizing photon loss in the reflected bosonic pulse (either from imperfect cavity coupling or atomic state decay). Imperfect phases on the reflected coherent pulses leads to incorrect displacement of the GKP states, resulting in effective Pauli errors or states lying outside the GKP logical basis. Photon loss in the cavity-photon interface reduces the overall strength of the reflected pulses and results in an effective dephasing noise on the final state. Considering these realistic effects, given the initial memory state initialization as in Eq. 7, the final state can most generally be described by,
$\begin{split}\rho_{\bm{M}G}\approx\frac{1}{d}\sum_{\bm{m},\bm{m}^{\prime}=0}^{d-1}g_{\bm{m},\bm{m}^{\prime}}\ket{\bm{m}}\!\!\bra{\bm{m}^{\prime}}_{\bm{M}}D(\bm{\beta}_{\bm{m}^{\prime}})\rho_{G}D^{\dagger}(\bm{\beta}_{\bm{m}^{\prime}}),\end{split}$ (11)
where corresponding to the memory basis state |$\ket{\bm{m}}_{\bm{M}}$ the realistic displacement imparted is
@@
$g_{\bm{m},\bm{m}^{\prime}}=\delta_{\bm{m},\bm{m}^{\prime}}+(1-\delta_{\bm{m},\bm{m}^{\prime}})|\lambda_{\bm{m},\bm{m}^{\prime}}|^{2},$ (13)
with $|\lambda_{\bm{m},\bm{m}^{\prime}}|^{2}$ representing the effective dephasing factors due to photon loss and $\delta_{\bm{m},\bm{m}^{\prime}}$ is the Kronecker delta symbol. For $C\gg 1$ and $|\omega-\omega_{c}|\rightarrow 0$, we may achieve the approximate relation $-r^{(0)*}(\omega)\approx r^{(1)*}(\omega)=1$ over the spectral bandwidth of the pulses (see Methods for details). Any additional sources of loss can be incorporated by including sub-unity multiplicative factors for $|\lambda_{\bm{m},\bm{m}^{\prime}}|^{2}$. The efficacy of a non-ideal protocol implementation for $N=1;d=2$ (i.e. CNOT/CPHASE gate between a GKP qubit and a single memory) is discussed in detail in Ref. [2].
@@
Figure 3: Layout of high-rate quantum links utilizing GKP qudit basis entanglement swap operation, where each party performs the CSUM gate on local memory registers $(\bm{M}_{1},\bm{M}_{2})$ and GKP qudit states $(G_{1},G_{2})$ before transmission to the midpoint of a link (of transmissivity $\eta$). The dual-homodyne measurements $\hat{q}_{1},\hat{p}_{2}$ herald the generation of an entangled state of the memory registers $(\tilde{\rho}_{\bm{M}_{1},\bm{M}_{2}})$.
@@
The effect of loss (with implicit pre-amplification of the state [7] or post-processing [5] of measurement outcomes) and finite-squeezing [19] on the GKP qudit states can be effectively modeled by a Gaussian random displacement channel on the ideal qudit states [5, 20, 7]. Given an initial state described by the density operator $\rho_{G}$, the random displacement channel $\mathcal{N}_{B_{2}}[\sigma^{2}]$ modifies the state as
@@
Lossless transmission i.e. $\eta=1$ yields a final state, |$\ket{\Psi_{k^{\prime},l^{\prime}}}_{\bm{M}_{1},\bm{M}_{2}}\equiv W_{1}^{(0,d-x_{L})}W_{2}^{(d-y_{L},0)}\ket{\Psi_{0,0}}_{\bm{M}_{1},\bm{M}_{2}}$, a different qudit Bell state; since both $x_{L}$ and $y_{L}$ have $d$ possible values, the heralded state is one of $d^{2}$ qudit Bell states (see Methods for notation related to qudit gates).
Introducing any loss (even with ideal GKP states) yields a final state (say $\rho_{\bm{M}_{1},\bm{M}_{2}}$) that is mixed. Assuming that the parties have access to finitely-squeezed GKP states of peak variance $\sigma_{i}^{2}$, transmission over a pure loss channel of transmissivity $\eta$ can be treated as an effective transformation of the variance. We consider the variance transformation under two protocol choices - pre-amplification of the transmitted state (equivalent to applying an amplification channel of gain $G=1/\eta$ before transmission), or, classical post-processing of the dual-homodyne measurement outcomes by a classical computer, introduced as ‘classical computer assisted amplification’ (CC-amplification) in Ref. [5]. The transformations can be summarized as
[…]
It is straightforward to note that CC-amplification adds a smaller value of noise for all values of loss $\eta>0.5$.
Given a state variance of $\sigma^{2}$, the probability of making $k$-shift errors in the logical measurement for sq-GKP states (along either the Pauli $X$ or $Z$ basis) is given by [7, 21],
$\begin{split}P^{(\mathrm{sq})}(X^{k},\sigma^{2})&\equiv P^{(\mathrm{sq})}(Z^{k},\sigma^{2})\\ &=\sum_{j\in\mathbb{Z}}\frac{1}{2}\Biggl{[}\mathrm{erf}\left(\sqrt{\frac{2\pi}{d}}\frac{jd+k+1/2}{\sigma}\right)\\ &\qquad\qquad\quad-\mathrm{erf}\left(\sqrt{\frac{2\pi}{d}}\frac{jd+k-1/2}{\sigma}\right)\Biggr{]},\end{split}$ (19)
where $k\in\{0,1,2,\ldots,d-1\}$. $k=0$ denotes the probability of incurring no shifts; the summation over the index $j$ accounts for shifts of magnitude greater than one lattice spacing of $\sqrt{2\pi d}$. Additionally, instead of exactly evaluating the effects of the random displacement channel, we consider a ‘twirled’ noise model. This lets us approximate $\rho_{\bm{M}_{1},\bm{M}_{2}}$ by the state,
$\tilde{\rho}_{\bm{M}_{1}\bm{M}_{2}}=\sum_{k_{1},k_{2}=0}^{d-1}P_{\mathrm{shift}}^{(k_{1},k_{2})}(\sigma^{2})\,\ket{\Psi_{k^{\prime},l^{\prime}}}\!\!\bra{\Psi_{k^{\prime},l^{\prime}}}_{\bm{M}_{1}\bm{M}_{2}},$ (20)
where $P_{\mathrm{shift}}^{(k_{1},k_{2})}(\sigma^{2})=P^{(\mathrm{sq})}(X^{k_{1}},\sigma^{2})\cdot P^{(\mathrm{sq})}(Z^{k_{2}},\sigma^{2})$. Further analysis in this article is limited to $\tilde{\rho}_{\bm{M}_{1}\bm{M}_{2}}$; we have elaborated on why this model of the output state serves as a mathematically simple and conservative estimate (in terms of the achievable entanglement generation rate) for the final state in the Methods. An achievable lower bound to the distillable entanglement for $\tilde{\rho}_{\bm{M}_{1}\bm{M}_{2}}$ is given by the hashing bound [22],